Assume that GDP is $6,000, personal disposable income is $5,100, and the government bud


get deficit is $200. Consumption is $3,800, and the trade deficit is $100.


a. How large is saving (S)? b. How large is investment (1)?


c. How large is government spending (G)?


1
Expert's answer
2021-08-10T17:44:49-0400

Y = Rs. 6000, Yd = Rs. 5100, BD = 200, C = Rs. 3800, NX = -1000.

a) How large is the saving

Saving S = Yd - C = 5100 - 3800 = Rs. 1300


Y = C + I + G + NX,


b) How large is the investment

Investment I = Y - C - G - NX = 6000 - 3800 - 1100 + 1000 = Rs. 2100


c)How large is the government spending

Government spending G = T + BD, T = Y - Yd = 6000 - 5100 = 900, so G = 900 + 200 = Rs. 1100.